OpenStudy (anonymous):

could someone please check these? 1) x^2-6x+8=(x-4)(x-2) 2) x^2+x-6=(x+2)(x-3) 3) x^2-5x+4=(x-4)(x-1) 4) x^2+x-2=(x+2)(x-1) 5) x^2-2x-8=(x+4)(x-2) 6) x^2+x-12=(x-3)(x+4) 7) x^2-x-12=(x+3)(x-4) 8) x^2-3x-4=(x-4)(x+1) 9) x^2-3x+2=(x-2)(x-1) 10) x^2-8x+16=(x-4)(x-4) 11) x^2-2x-15=(x+3)(x-5) 12) x^2+5x+6=(x+3)(x+2) 13) x^2-5x+6=(x-3)(x-2) 14) x^2+x-20=(x-4)(x+5) 15) x^2-x-2=(x+1)(x-2) 16) x^2+4x-5=(x-1)(x+5) 17) x^2+2x-8=(x-2)(x+4) 18) x^2+3x-10=(x-2)(x+5)

OpenStudy (anonymous):

1)x^2-6x+8=(x-4)(x-2) correct 2)x^2+x-6=(x+3)(x-2) incorrect

OpenStudy (anonymous):

3)x^2-5x+4=(x-4)(x-1) correct

OpenStudy (anonymous):

4) x^2+x-2=(x+2)(x-1) correct

OpenStudy (anonymous):

so #2 is (x-2)(x+3)

OpenStudy (anonymous):

5)x^2-2x-8=(x+4)(x-2) incorrect

OpenStudy (anonymous):

No.2 is (x+3)(x-2)

OpenStudy (anonymous):

no.5 is (x+2)(x-4)

OpenStudy (anonymous):

try doing these like you would any normal sum: x+3 x-2 ------ then multiply each one as follows: -2 by +3 = -6 then -2 by x is -2x then x by +3 = 3x and x by x = x^2 then collecting these you get -2x +3x = x so the whole thing is x^2 +x -6

OpenStudy (anonymous):

You are correct with No.5 (x+4)(x-2)

OpenStudy (anonymous):

Do you need me to check them all, or do you think you can now check them yourself?

OpenStudy (anonymous):

i've got it from here thank you soooo much you're a life saver (my final is tomorrow) i would give you 5 more medals if i could!!

OpenStudy (anonymous):

No worries, as long as you're okay, good luck in the finals!